# Can Slow-thinking LLMs Reason Over Time? Empirical Studies in Time Series Forecasting

### v1

@@ A.1 Time Series Forecasting
Time series forecasting has been extensively studied over the past few decades cheng2025comprehensive . Early methods, such as ARIMA box1968some ; zhang2003time , employed an auto-regressive approach and forecasted using a moving average framework. While theoretical guarantees were established, these traditional methods often require data with ideal properties, which do not always align with real-world scenarios hewamalage2021recurrent . With the growing availability of data and advances in computing power, numerous deep learning-based models have emerged, typically following a sequence-to-sequence paradigm salinas2020deepar . Recurrent neural networks were among the first to capture temporal dependencies by summarizing past information in time series data wen2017multi . However, these architectures suffer from limitations such as a restricted receptive field and error accumulation due to recursive inference lai2018modeling , which undermine forecasting accuracy. To improve performance, advanced architectures have been developed to capture long-range dependencies zhou2021informer , including those with self-attention mechanisms and convolutional networks li2019enhancing ; liu2022scinet . Recent work has also integrated traditional techniques, like trend-seasonal decomposition and time-frequency transformations zhou2022fedformer ; liu2024generative , into neural networks to better capture time series data characteristics wen2020fast ; zhou2022film ; cheng2025convtimenet . Additionally, a recent study showed that even a simple linear network enhanced with decomposition can achieve competitive performance zeng2023transformers . Furthermore, slice-based methods have demonstrated superior accuracy in long-term forecasting tasks nie2022time ; zhang2023crossformer ; cheng2025instructime .
@@ A.2 Large Language Models for Time Series
In recent years, large language models (LLMs) have garnered significant attention for their remarkable ability to understand and generate human-like text kumar2024large ; yao2024survey . While traditionally applied to natural language processing tasks zhang2024large , these models have also shown great potential in time series analysis  jin2024position . The application of LLMs to time series can be broadly categorized into two primary paradigms: fine-tuning and prompt-based zero-shot learning jiang2024empowering . In fine-tuning, a pre-trained LLM is further trained on time series data specific to a given task chang2023llm4ts ; chang2024align , enabling it to learn task-specific patterns and improve performance  liu2025calf ; jin2023time . However, this approach often requires substantial labeled data and computational resources cheng2025comprehensive . On the other hand, prompt-based zero-shot methods leverage the model’s pre-existing knowledge gruver2023large , guiding its predictions through well-crafted prompts without the need for task-specific training liu2024lstprompt . Although zero-shot methods are more flexible and resource-efficient tang2025time , they may not always achieve the same level of performance as fine-tuned models, especially on specialized time series tasks wang2024tabletime . Both paradigms highlight the growing interest in applying LLMs to time series, although challenges remain in optimizing effectiveness for these tasks.
@@ A.3 Inference-Time Reasoning
Inference-Time Reasoning zeng2025revisiting refers to the strategy of enhancing a pre-trained model’s performance on complex reasoning tasks by allocating additional computational resources during the inference phase singhi2025solve ; yang2025towards . These resources may include extended inference time, more computational steps, or advanced reasoning strategies such as multi-path sampling and iterative refinement leehanchung . Unlike traditional training-time scaling, which focuses on improving a model’s capabilities during training through increased parameters, data, or computation parashar2025inference , inference-time reasoning utilizes a larger computational budget at inference to unlock deeper reasoning potential within the model—potential zhang2025and that might otherwise remain dormant during rapid inference. This approach improves performance without altering the model’s weights chen2024simple .
Several cutting-edge large language models demonstrate the effectiveness of inference time reasoning sui2025meta . For instance, DeepSeek-R1 guo2025deepseek can benefit from techniques such as iterative refinement or self-correction, utilizing additional computational resources during inference to enhance performance on complex tasks, leveraging its strengths in coding and reasoning mercer2025brief . Similarly, models such as Gemini-2.5-pro team2023gemini are expected to excel in sophisticated reasoning tasks when granted extended processing time or multi-step inference protocols enabled by inference-time reasoning. OpenAI’s models, including versions such as GPT-o1 hasei2025comparative , show significant improvements on challenging reasoning problems when employing techniques such as chain-of-thought prompting or step-by-step decomposition, both of which increase the computational budget during inference wei2022chain . These examples highlight the growing importance of inference-time reasoning in pushing the boundaries of AI reasoning by trading inference-time computation for higher accuracy muennighoff2025s1 .
@@ B.2 Baselines
  • Autoformer wu2021autoformer : Introduces a decomposition-based Transformer architecture with an auto-correlation mechanism to capture long-term dependencies in time series forecasting.
  • FEDformer zhou2022fedformer : Combines seasonal-trend decomposition with frequency domain analysis using Fourier transforms to enhance long-term time series forecasting.
  • PatchTST nie2022time : Segments time series into patches and applies a channel-independent Transformer to improve long-term forecasting accuracy.
  • iTransformer liu2023itransformer : Applies attention mechanisms on inverted dimensions, embedding time points as variate tokens to capture multivariate correlations in time series data.
  • TimeXer wang2024timexer : Integrates exogenous variables into Transformer-based time series forecasting, enhancing predictions by incorporating external information.
  • DLinear zeng2023transformers : A simple and efficient linear model that decomposes time series into trend and seasonal components, achieving competitive performance with reduced complexity.
  • GPT4TS zhou2023one : Utilizes pre-trained GPT-2 models for time series forecasting by feeding time series patches in a channel-independent manner.
  • Time-LLM jin2023time : Reprograms large language models for time series forecasting by converting time series data into textual prompts, enabling few-shot and zero-shot learning capabilities.
  • LLMTime gruver2023large : Employs large language models for zero-shot time series forecasting by encoding numerical data as text and generating predictions through text completions.
@@ B.5 Normalization Strategies
Normalization is a key preprocessing step in time series forecasting. We compare two strategies: reversible instance normalization (RevIN) kim2021reversible and Z-score normalization, which differ in how they compute and apply normalization statistics.

### v2

@@ A.1 Time Series Forecasting
Time series forecasting has been extensively studied over the past few decades [11]. Early methods, such as ARIMA [3, 76], employed an auto-regressive approach and forecasted using a moving average framework. While theoretical guarantees were established, these traditional methods often require data with ideal properties, which do not always align with real-world scenarios [22]. With the growing availability of data and advances in computing power, numerous deep learning-based models have emerged, typically following a sequence-to-sequence paradigm [55]. Recurrent neural networks were among the first to capture temporal dependencies by summarizing past information in time series data [67]. However, these architectures suffer from limitations such as a restricted receptive field and error accumulation due to recursive inference [33], which undermine forecasting accuracy. To improve performance, advanced architectures have been developed to capture long-range dependencies [82], including those with self-attention mechanisms and convolutional networks [37, 43]. Recent work has also integrated traditional techniques, like trend-seasonal decomposition and time-frequency transformations [84, 47], into neural networks to better capture time series data characteristics [66, 83, 12]. Additionally, a recent study showed that even a simple linear network enhanced with decomposition can achieve competitive performance [73]. Furthermore, slice-based methods have demonstrated superior accuracy in long-term forecasting tasks [50, 81, 10].
@@ A.2 Large Language Models for Time Series
In recent years, large language models (LLMs) have garnered significant attention for their remarkable ability to understand and generate human-like text [32, 71]. While traditionally applied to natural language processing tasks [80], these models have also shown great potential in time series analysis  [28]. The application of LLMs to time series can be broadly categorized into two primary paradigms: fine-tuning and prompt-based zero-shot learning [26]. In fine-tuning, a pre-trained LLM is further trained on time series data specific to a given task [4, 5], enabling it to learn task-specific patterns and improve performance  [44, 27]. However, this approach often requires substantial labeled data and computational resources [11]. On the other hand, prompt-based zero-shot methods leverage the model’s pre-existing knowledge [19], guiding its predictions through well-crafted prompts without the need for task-specific training [41]. Although zero-shot methods are more flexible and resource-efficient [59], they may not always achieve the same level of performance as fine-tuned models, especially on specialized time series tasks [63]. Both paradigms highlight the growing interest in applying LLMs to time series, although challenges remain in optimizing effectiveness for these tasks.
@@ A.3 Inference-Time Reasoning
Inference-Time Reasoning [75] refers to the strategy of enhancing a pre-trained model’s performance on complex reasoning tasks by allocating additional computational resources during the inference phase [56, 70]. These resources may include extended inference time, more computational steps, or advanced reasoning strategies such as multi-path sampling and iterative refinement [35]. Unlike traditional training-time scaling, which focuses on improving a model’s capabilities during training through increased parameters, data, or computation [51], inference-time reasoning utilizes a larger computational budget at inference to unlock deeper reasoning potential within the model—potential [77] that might otherwise remain dormant during rapid inference. This approach improves performance without altering the model’s weights [9].
Several cutting-edge large language models demonstrate the effectiveness of inference time reasoning [57]. For instance, DeepSeek-R1 [20] can benefit from techniques such as iterative refinement or self-correction, utilizing additional computational resources during inference to enhance performance on complex tasks, leveraging its strengths in coding and reasoning [48]. Similarly, models such as Gemini-2.5-pro [61] are expected to excel in sophisticated reasoning tasks when granted extended processing time or multi-step inference protocols enabled by inference-time reasoning. OpenAI’s models, including versions such as GPT-o1 [21], show significant improvements on challenging reasoning problems when employing techniques such as chain-of-thought prompting or step-by-step decomposition, both of which increase the computational budget during inference [65]. These examples highlight the growing importance of inference-time reasoning in pushing the boundaries of AI reasoning by trading inference-time computation for higher accuracy [49].
@@ B.2 Baselines
  • Autoformer [68]: Introduces a decomposition-based Transformer architecture with an auto-correlation mechanism to capture long-term dependencies in time series forecasting.
  • FEDformer [84]: Combines seasonal-trend decomposition with frequency domain analysis using Fourier transforms to enhance long-term time series forecasting.
  • PatchTST [50]: Segments time series into patches and applies a channel-independent Transformer to improve long-term forecasting accuracy.
  • iTransformer [45]: Applies attention mechanisms on inverted dimensions, embedding time points as variate tokens to capture multivariate correlations in time series data.
  • TimeXer [64]: Integrates exogenous variables into Transformer-based time series forecasting, enhancing predictions by incorporating external information.
  • DLinear [73]: A simple and efficient linear model that decomposes time series into trend and seasonal components, achieving competitive performance with reduced complexity.
  • GPT4TS [85]: Utilizes pre-trained GPT-2 models for time series forecasting by feeding time series patches in a channel-independent manner.
  • Time-LLM [27]: Reprograms large language models for time series forecasting by converting time series data into textual prompts, enabling few-shot and zero-shot learning capabilities.
  • LLMTime [19]: Employs large language models for zero-shot time series forecasting by encoding numerical data as text and generating predictions through text completions.
@@ B.5 Normalization Strategies
Normalization is a key preprocessing step in time series forecasting. We compare two strategies: reversible instance normalization (RevIN) [30] and Z-score normalization, which differ in how they compute and apply normalization statistics.